40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 22 21:30:06 2008, rana, Configuration, PSL, bad FSS DAQ.png
    Reply  Mon Sep 22 22:24:19 2008, rana, Configuration, PSL, bad FSS 
Message ID: 980     Entry time: Mon Sep 22 21:30:06 2008     Reply to this: 982
Author: rana 
Type: Configuration 
Category: PSL 
Subject: bad FSS 
The MC refl power was going up and the FSS PC drive was so large that I had to turn up the FSS
common gain from 1.5 dB to 10.5 dB to get it to be better. Attached are the before (REF) and
after plots of frequency noise. Is the FSS gain really supposed to be 1.5 dB?? How did we gain
so many dB's of optical gain? Is there a loop measurement from after Peter's oscillator change?
Attachment 1: DAQ.png  66 kB  | Hide | Hide all
DAQ.png
ELOG V3.1.3-